OpenStudy (anonymous):

Inverse trig function: -5cot^-1 [x/6] find dy/dx. I have alot of problems like this and I need help with one so I can do the rest

OpenStudy (anonymous):

The formula for the derivative of the inverse cotangent is \[\frac{ d }{dx}(arccot)x = -\frac{ 1 }{ 1 + x^2 }\] For this problem after you take the derivative should look something like this after applying chain rule and prior to simplification\[-5[-\frac{ 1 }{ 1 + (\frac{ x }{ 6})^2}]\frac{ 1 }{ 6}\] After some cleaning up would look like \[\frac{ 5 }{ 6 }(1 + (\frac{ x }{ 6 })^2)^{-1}\]

OpenStudy (anonymous):

Is that the final answer?

OpenStudy (anonymous):

If you are taking d/dx. Is this for calc 1?

OpenStudy (anonymous):

Since this is already solved for y, dy/dx is in essence the same as d/dx. And is the final although could choose other ways to simplify.